What is the x-intercept with the vertex(1,-17)and y-intercept(0,-16)
Answer by lwsshak3(11628)

What is the x-intercept with the vertex(1,-17)and y-intercept(0,-16)
**
With coordinates of vertex given,
equation can be written as follows:
y=(x-1)^2-17
To find the x-intercepts, set y=0 and solve for x
0=(x-1)^2-17
(x-1)^2=17 take sqrt of both sides
x-1=±√17≈±4.123
x=1+√17≈1+4.123≈5.123
or
x=1-√17≈1-4.123≈-3.123
note: y-intercept (0,-16) not needed to find x-intercepts